OpenStudy (anonymous):

HELPPP!!!!!ARE THE ANSWERS BELOW CORRECT? 1. Solve the equation. Round the answer to the nearest tenth. 3.6x = 5.5 (Points : 1) 15.3 1.6 1.5 1.4

OpenStudy (anonymous):

2. Solve the equation. Round the answer to the nearest tenth. 1.18x + 4.34 = 8.67 (Points : 1) 11.0 3.8 3.7 3.6 3. Estimate the solution to the equation by first rounding the constants and coefficients to the nearest integer. 3.1x + 5.2 = 9.9 (Points : 1) 1.3 1.5 1.6 1.7 4. Estimate the solution to the equation by first rounding the constants and coefficients to the nearest integer. 4.12(x – 7.89) = 18.12 (Points : 1) 12.2 12.4 12.5 12.6

OpenStudy (anonymous):

1.A 2.B 3.D 4.C Are these answers correct?